# A Careful Examination of Large Language Model Performance on Grade School Arithmetic

### 03 May 2024 v3

@@ 6 Discussion
We create GSM1k, a novel dataset designed to measure LLM overfitting on GSM8k. When benchmarking leading open- and closed-source models, we find substantial evidence that many models have been contaminated by benchmark data, with models showing performance drops of up to 13% accuracy. Additionally, we find that several families of models, most notably the Mistral and Phi families, show consistent overfitting across almost all model sizes and versions. An extended analysis reveals a positive relationship between a model’s likelihood of generating data points in GSM8k and its performance difference between GSM8k and GSM1k, suggesting evidence of data contamination as one of the underlying causes. Nevertheless, we find that frontier models exhibit little to no evidence of overfitting and that many models, even the most heavily overfit families, show strong signs of generalizable mathematical reasoning.
@@ 7 Acknowledgements
We would like to thank Dan Hendrycks, Adi Ganesh, Akilesh Praveen, Andrea Jaba, Charlotte Zhuang, Will Zhou, Celia Chen and Kamilė Lukošiūtė for their helpful comments and suggestions.
Appendix A Annotator Instructions
@@
Appendix B N-shot Prompt (examples selected randomly from GSM8k train)
@@
Appendix C Results with An Alternative Prompt
@@
We report our results in Table 1. On average, we find that this prompt causes the gap between GSM8k and GSM1k to decrease by about 1% across all models. However, for some models (e.g. Mixtral-8x22B-v0.1), this reduces the amount of observed overfitting dramatically. While the exact cause of this difference is impossible to know, especially without access to model details such as their training set, our hypothesis is that prompting a model with GSM8k is more likely to activate the “memorization” portion of a model than if it is prompted by non-GSM8k grade school math problems.
Appendix D Results Table
We report our full results in Table 1. Models are sorted by the difference in performance between GSM8k and GSM1k. Because all models are evaluated using the standard LM Evaluation Harness prompt and evaluation format, model performance on GSM8k may not match reported benchmark numbers. In particular, answers that do not match the 5-shot example format are marked incorrect even if they are otherwise “correct.” Our focus is primarily on the difference between GSM8k and GSM1k performance, holding evaluation setting constant. The Z-score and p-value are calculated for a two-tailed two proportion Z-test. Alternative prompt results are also included. For details, see Appendix C.
@@ Appendix D Results Table
math-shepherd-mistral-7b-rl 0.134 0.745 0.611 7.322 0.000
Xwin-Math-13B-V1.0 0.101 0.631 0.529 5.196 0.000
Xwin-Math-7B-V1.0 0.101 0.529 0.428 5.131 0.000
Mixtral-8x22B-Instruct-v0.1 0.099 0.859 0.760 6.403 0.000
Phi-3-mini-4k-instruct 0.098 0.782 0.684 5.600 0.000
deepseek-llm-67b-base 0.093 0.615 0.522 4.771 0.000
Mixtral-8x22B-v0.1 0.092 0.770 0.677 5.257 0.000
gemma-7b 0.077 0.541 0.464 3.879 0.000
phi-2 0.074 0.569 0.495 3.766 0.000
Yi-6B-Chat 0.073 0.408 0.334 3.851 0.000
mistral-small-latest 0.072 0.790 0.718 4.218 0.000
Yi-34B-Chat 0.071 0.641 0.569 3.683 0.000
Smaug-34B-v0.1 0.069 0.757 0.688 3.931 0.000
gemma-1.1-7b-it 0.069 0.522 0.453 3.485 0.000
falcon-180B-chat 0.065 0.610 0.545 3.320 0.001
codegemma-7b 0.063 0.479 0.416 3.217 0.001
CodeLlama-70b-Python-hf 0.063 0.462 0.399 3.198 0.001
command 0.062 0.445 0.383 3.221 0.001
@@
Smaug-2-72B 0.061 0.799 0.738 3.697 0.000
CodeLlama-70b-hf 0.059 0.486 0.427 2.989 0.003
Phi-3-mini-128k-instruct 0.058 0.741 0.683 3.264 0.001
CodeLlama-34b-Instruct-hf 0.058 0.415 0.357 3.013 0.003
Meta-Llama-3-8B 0.057 0.499 0.442 2.865 0.004
codegemma-7b-it 0.055 0.526 0.471 2.784 0.005
phi-1 0.048 0.324 0.276 2.683 0.007
Mixtral-8x7B-v0.1 0.048 0.578 0.530 2.450 0.014
Phind-CodeLlama-34B-v2 0.048 0.416 0.368 2.460 0.014
Mixtral-8x7B-Instruct-v0.1 0.046 0.641 0.594 2.411 0.016
gemma-7b-it 0.044 0.335 0.291 2.397 0.017
mistral-medium-latest 0.038 0.790 0.752 2.292 0.022
vicuna-33b-v1.3 0.038 0.379 0.341 2.020 0.043
CodeLlama-13b-Python-hf 0.030 0.214 0.183 1.942 0.052
OpenMath-Llama-2-70b-hf 0.029 0.171 0.142 2.072 0.038
Mistral-7B-v0.1 0.029 0.381 0.353 1.500 0.134
CodeLlama-34b-hf 0.028 0.350 0.322 1.497 0.134
CodeLlama-13b-Instruct-hf 0.027 0.277 0.250 1.560 0.119
Mistral-7B-Instruct-v0.2 0.027 0.426 0.399 1.383 0.167
CodeLlama-70b-Instruct-hf 0.025 0.497 0.471 1.287 0.198
Llama-2-7b-hf 0.023 0.141 0.118 1.766 0.077
Meta-Llama-3-70B 0.022 0.811 0.789 1.371 0.170
Llama-2-70b-hf 0.021 0.538 0.516 1.092 0.275
Meta-Llama-3-70B-Instruct 0.020 0.896 0.876 1.607 0.108
Mistral-7B-Instruct-v0.1 0.019 0.335 0.316 1.033 0.302
gemini-1.5-pro-preview-0409 0.018 0.897 0.879 1.423 0.155
pythia-12b 0.016 0.036 0.021 2.362 0.018
CodeLlama-13b-hf 0.015 0.227 0.212 0.899 0.369
CodeLlama-34b-Python-hf 0.013 0.306 0.293 0.746 0.456
gemma-2b 0.013 0.176 0.163 0.911 0.362
dbrx-base 0.013 0.727 0.715 0.715 0.474
deepseek-coder-33b-instruct 0.010 0.419 0.409 0.538 0.591
CodeLlama-7b-Python-hf 0.008 0.127 0.120 0.629 0.529
gpt-3.5-turbo 0.007 0.760 0.753 0.405 0.685
gpt2-xl 0.004 0.021 0.017 0.665 0.506
gpt-neox-20b 0.003 0.059 0.056 0.253 0.800
gemini-pro 0.002 0.792 0.789 0.119 0.905
CodeLlama-7b-hf 0.002 0.127 0.125 0.135 0.893
CodeLlama-7b-Instruct-hf -0.000 0.186 0.186 0.010 0.992
mistral-large-latest -0.000 0.853 0.853 0.009 0.993
gpt-4-turbo -0.000 0.898 0.898 0.004 0.997
claude-3-haiku-20240307 -0.000 0.785 0.785 -0.007 0.994
gemma-2b-it -0.001 0.111 0.112 -0.044 0.965
Llama-2-13b-hf -0.004 0.230 0.235 -0.283 0.777
claude-2.1 -0.007 0.887 0.894 -0.532 0.595
gpt-4 -0.007 0.911 0.918 -0.638 0.523
gemma-1.1-2b-it -0.008 0.104 0.112 -0.664 0.506
deepseek-math-7b-rl -0.022 0.185 0.207 -1.418 0.156
claude-3-opus-20240229 -0.023 0.802 0.825 -1.474 0.141
claude-3-sonnet-20240229 -0.024 0.719 0.744 -1.355 0.175
@@
math-shepherd-mistral-7b-rl 0.138 0.782 0.645 7.726 0.000
deepseek-math-7b-rl 0.108 0.754 0.646 5.979 0.000
Xwin-Math-13B-V1.0 0.076 0.643 0.567 3.966 0.000
Phi-3-mini-128k-instruct 0.065 0.817 0.752 4.029 0.000
gemma-1.1-7b-it 0.065 0.493 0.427 3.334 0.001
Yi-34B-Chat 0.062 0.647 0.584 3.266 0.001
deepseek-llm-67b-base 0.061 0.656 0.594 3.215 0.001
Yi-6B-Chat 0.061 0.423 0.362 3.187 0.001
CodeLlama-34b-Python-hf 0.059 0.340 0.281 3.216 0.001
Xwin-Math-7B-V1.0 0.058 0.513 0.456 2.903 0.004
codegemma-7b 0.057 0.516 0.459 2.896 0.004
Smaug-2-72B 0.057 0.812 0.755 3.497 0.000
codegemma-7b-it 0.056 0.516 0.459 2.855 0.004
CodeLlama-70b-Python-hf 0.054 0.493 0.439 2.722 0.006
command 0.049 0.446 0.396 2.513 0.012
phi-1 0.048 0.324 0.276 2.637 0.008
Smaug-34B-v0.1 0.046 0.745 0.700 2.607 0.009
Mixtral-8x22B-Instruct-v0.1 0.042 0.885 0.843 3.077 0.002
CodeLlama-70b-Instruct-hf 0.041 0.519 0.477 2.117 0.034
phi-2 0.041 0.554 0.513 2.103 0.035
CodeLlama-70b-hf 0.039 0.500 0.460 2.009 0.045
OpenMath-Llama-2-70b-hf 0.039 0.165 0.126 2.796 0.005
Phi-3-mini-4k-instruct 0.039 0.801 0.763 2.392 0.017
dbrx-base 0.033 0.718 0.685 1.837 0.066
Mistral-7B-Instruct-v0.2 0.033 0.436 0.403 1.680 0.093
mistral-small-latest 0.031 0.782 0.751 1.871 0.061
Mixtral-8x7B-Instruct-v0.1 0.029 0.686 0.657 1.540 0.124
deepseek-coder-33b-instruct 0.028 0.421 0.392 1.484 0.138
CodeLlama-13b-Python-hf 0.027 0.221 0.194 1.683 0.092
gemma-7b 0.026 0.544 0.518 1.318 0.188
Phind-CodeLlama-34B-v2 0.026 0.403 0.376 1.381 0.167
Meta-Llama-3-8B-Instruct 0.026 0.772 0.746 1.505 0.132
falcon-180B-chat 0.025 0.622 0.597 1.292 0.196
mistral-medium-latest 0.025 0.789 0.764 1.488 0.137
pythia-12b 0.023 0.050 0.027 3.111 0.002
vicuna-33b-v1.3 0.023 0.440 0.417 1.171 0.241
CodeLlama-7b-Instruct-hf 0.023 0.187 0.164 1.548 0.122
CodeLlama-7b-hf 0.021 0.116 0.095 1.781 0.075
Mixtral-8x22B-v0.1 0.019 0.810 0.791 1.173 0.241
CodeLlama-34b-hf 0.019 0.324 0.305 1.033 0.302
Meta-Llama-3-8B 0.018 0.547 0.529 0.904 0.366
gemini-1.5-pro-preview-0409 0.016 0.908 0.892 1.376 0.169
Llama-2-7b-hf 0.015 0.143 0.127 1.136 0.256
Meta-Llama-3-70B-Instruct 0.015 0.899 0.885 1.173 0.241
gpt-4 0.014 0.919 0.904 1.259 0.208
gemma-2b 0.013 0.191 0.178 0.879 0.380
CodeLlama-13b-Instruct-hf 0.012 0.287 0.276 0.684 0.494
gemma-7b-it 0.008 0.248 0.240 0.467 0.641
CodeLlama-34b-Instruct-hf 0.007 0.400 0.392 0.391 0.696
Mistral-7B-v0.1 0.007 0.422 0.415 0.325 0.745
Mixtral-8x7B-v0.1 0.006 0.607 0.601 0.336 0.737
Llama-2-70b-hf 0.006 0.572 0.566 0.309 0.757
gpt-neox-20b 0.005 0.076 0.070 0.527 0.598
gpt2-xl 0.004 0.022 0.018 0.645 0.519
CodeLlama-7b-Python-hf 0.003 0.123 0.120 0.281 0.779
CodeLlama-13b-hf 0.003 0.218 0.215 0.147 0.883
Llama-2-13b-hf 0.002 0.281 0.279 0.117 0.907
claude-2.1 0.001 0.836 0.836 0.016 0.987
gemini-pro 0.000 0.688 0.687 0.024 0.981
claude-3-haiku-20240307 0.000 0.792 0.791 0.019 0.985
gpt-4-turbo -0.002 0.847 0.849 -0.137 0.891
mistral-large-latest -0.003 0.854 0.857 -0.225 0.822
gemma-1.1-2b-it -0.006 0.084 0.090 -0.490 0.624
gemma-2b-it -0.007 0.099 0.106 -0.589 0.556
Mistral-7B-Instruct-v0.1 -0.007 0.347 0.354 -0.380 0.704
claude-3-sonnet-20240229 -0.011 0.713 0.724 -0.596 0.551
gpt-3.5-turbo -0.016 0.742 0.759 -0.946 0.344
claude-3-opus-20240229 -0.026 0.830 0.857 -1.855 0.064
Meta-Llama-3-70B -0.035 0.807 0.841 -2.322 0.020
Appendix E 50 Examples from GSM1k
@@
4 Colin is playing dice with his friend Eoin and needs some help keeping track of his score. He begins with 5 points and wins 6 points in the first round. In the second round, he won twice as many points as he won in the first round. In the third round, he had a fantastic roll and was able to triple his current total point count! How many points did Colin end the game with? 69
5 Bradley and his friends enjoy playing marbles. They possess a box of marbles containing 12 red balls, 15 yellow balls, and 18 green balls. How many additional red balls do they require to double the number of red balls compared to the combined number of yellow and green balls? 54
@@
10 Emily is applying to 6 different colleges. ½ of the colleges have an application fee of \$60, and the other half have an application fee of \$90. She must also pay \$15 per transcript to send them to each college. Her parents offer to help pay for half of the total costs. How many dollars does she have to pay? 270
@@
22 Carolina is trying to qualify for a car loan. The lender tells her she must meet a debt-to-income ratio of 1:4. Her current debts are \$900 in rent, \$200 in utilities, and another \$300 in miscellaneous expenses per month. Her current monthly salary is \$4000. How much more money, in dollars, will she need to cut out from her current debts per month to meet the DTI requirements? 400
@@
24 Frank sneaks out before his break at 3:20 pm and gets back at 4:05. If his break was only supposed to be half an hour, for how much longer did Frank sneak out (in minutes)? 15
25 Janet wants to listen to 20 music albums by the end of the week. If its Thursday and she just finished album number twelve and she has to finish them by Saturday, how many albums would she have to listen to per day? 4
@@
27 Gray has \$126 to spend on lunches for the week. On Monday, he spent \$16 on a carne asada burrito and a soda. On Saturday, he will spend \$30 eating out with friends. If he spends the same amount of money on food for the other 5 days of the week, what will be his average daily spending on food over these 5 days? 16
@@
47 Johnny has 8 green balls. He has five fewer than twice that number in red balls. How many total balls does Johnny have? 19
@@
Appendix F Additional Plots From Log-Likelihood Experiments
Figure 10: Log-likelihood of models on GSM1k (ours). As expected, we observe almost no correlation between the model’s probability of generating GSM1k and its level of overfit. This is because GSM1k is newly created and not on the internet.
Figure 11: Difference between log-likelihood of GSM1k (ours) and GSM8k.
Appendix G Bar Chart of Performance Gaps Between GSM8k and GSM1k Across All Model Accuracies

### 22 Nov 2024 v4

@@ 6 Discussion
We create GSM1k, a novel dataset designed to measure LLM overfitting on GSM8k. When benchmarking leading open- and closed-source models, we find substantial evidence that many models have been contaminated by benchmark data, with models showing performance drops of up to 8% accuracy. Additionally, we find that several model families show consistent overfitting across almost all model sizes and versions. An extended analysis reveals a positive relationship between a model’s likelihood of generating data points in GSM8k and its performance difference between GSM8k and GSM1k, suggesting evidence of data contamination as one of the underlying causes. Nevertheless, we find that frontier models exhibit little to no evidence of overfitting and that many models, even the most heavily overfit families, show strong signs of generalizable mathematical reasoning.
@@ 7 Acknowledgements
We would like to thank Dan Hendrycks, Adi Ganesh, Akilesh Praveen, Andrea Jaba, Will Zhou, Celia Chen, Apaar Shanker and Kamilė Lukošiūtė for their helpful comments and suggestions.
Appendix A Checklist
  1. 1.
    For all authors…
    1. (a)
      Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope? [Yes]
    2. (b)
      Did you describe the limitations of your work? [Yes]
    3. (c)
      Did you discuss any potential negative societal impacts of your work? [N/A]
    4. (d)
      Have you read the ethics review guidelines and ensured that your paper conforms to them? [Yes]
  2. 2.
    If you are including theoretical results…
    1. (a)
      Did you state the full set of assumptions of all theoretical results? [N/A]
    2. (b)
      Did you include complete proofs of all theoretical results? [N/A]
  3. 3.
    If you ran experiments (e.g. for benchmarks)…
    1. (a)
      Did you include the code, data, and instructions needed to reproduce the main experimental results (either in the supplemental material or as a URL)? [Yes] included in supplement, will open source shortly
    2. (b)
      Did you specify all the training details (e.g., data splits, hyperparameters, how they were chosen)? [N/A]
    3. (c)
      Did you report error bars (e.g., with respect to the random seed after running experiments multiple times)? [Yes] Appendix F includes standard errors and statistical significance.
    4. (d)
      Did you include the total amount of compute and the type of resources used (e.g., type of GPUs, internal cluster, or cloud provider)? [Yes] We run all models on a cluster with 8 x A100 nodes. Most models complete evaluation within a few minutes.
  4. 4.
    If you are using existing assets (e.g., code, data, models) or curating/releasing new assets…
    1. (a)
      If your work uses existing assets, did you cite the creators? [N/A]
    2. (b)
      Did you mention the license of the assets? [N/A]
    3. (c)
      Did you include any new assets either in the supplemental material or as a URL? [Yes] GSM1k is yet unreleased, with a future release date. We provide the full dataset here. We ask that reviewers refrain from sharing this dataset publicly.
    4. (d)
      Did you discuss whether and how consent was obtained from people whose data you’re using/curating? [Yes] All problems created by annotators hired by Scale AI.
    5. (e)
      Did you discuss whether the data you are using/curating contains personally identifiable information or offensive content? [N/A]
  5. 5.
    If you used crowdsourcing or conducted research with human subjects…
    1. (a)
      Did you include the full text of instructions given to participants and screenshots, if applicable? [Yes] See Appendix C
    2. (b)
      Did you describe any potential participant risks, with links to Institutional Review Board (IRB) approvals, if applicable? [N/A]
    3. (c)
      Did you include the estimated hourly wage paid to participants and the total amount spent on participant compensation? [Yes] Annotators were paid 20-25 / hour, depending on performance, experience, and bonus incentives. In total, Scale paid out around 180K to human annotators to create this benchmark.
Appendix B Dataset Documentation
  1. 1.
    Construction. GSM1k is a dataset of 1205 questions requiring elementary mathematical reasoning to solve. All problems are intended to be solvable using only the four basic arithmetic operators.
  2. 2.
    Creation. GSM1k was created using human annotation from scratch without any usage of LLMs. Human annotators were hired by Scale AI and paid between 20 and 25 dollars per hour. All annotators were based in the United States. In total, this dataset paid out around \$180,000 dollars to human annotators, including costs resulting from problem creation and solving, quality assurance checks, as well as experiments done to compare the difficulty distribution with GSM8k.
  3. 3.
    Intent. This dataset is intended to be used as a held-out version of GSM8k to measure data contamination. As such, it largely mimics the format and style of GSM8k. All answers are a non-negative integer.
  4. 4.
    Release. Our dataset is not published at this time, to prevent risk of data contamination in future models. We will release Croissant metadata when the dataset is public, with the conditions described in the main paper.
  5. 5.
    Liability. The authors bear all responsibility in case of violation of rights. Due to Scale AI commissioning the construction of this dataset from scratch primarily for this purpose of this paper, we do not anticipate any copyright or other issues. The dataset (yet unreleased) will be released with the MIT license.
  6. 6.
    Preservation. We plan to release the full dataset on Github as well as HuggingFace so it remains publicly accessible to anyone who wishes to use it. The formatting will be 1205 rows with a question and answer column.
Appendix C Annotator Instructions
@@
Appendix D N-shot Prompt (examples selected randomly from GSM8k train)
@@
Appendix E Results with an Alternative Prompt
@@
We report our results in Table 1. While there is significant variance based on prompt, the general trend of which model families are overfit is similar.
Figure 7: Gap in accuracy between GSM8k and GSM1k for models that score above 70% on GSM8k.
Appendix F Results Table
We report our full results in Table 1. Models are sorted by the difference in performance between GSM8k and GSM1k. Because all models are evaluated using the standard LM Evaluation Harness prompt and evaluation format, model performance on GSM8k may not match reported benchmark numbers. In particular, answers that do not match the 5-shot example format are marked incorrect even if they are otherwise “correct.” Our focus is primarily on the difference between GSM8k and GSM1k performance, holding evaluation setting constant. The Z-score and p-value are calculated for a two-tailed two proportion Z-test. Alternative prompt results are also included. For details, see Appendix E.
@@ Appendix F Results Table
Yi-6B-Chat 0.080 0.437 0.357 4.135 0.000
math-shepherd-mistral-7b-rl 0.072 0.826 0.754 4.488 0.000
command 0.065 0.447 0.383 3.336 0.000
Xwin-Math-13B-V1.0 0.064 0.660 0.596 3.334 0.000
phi-2 0.063 0.566 0.504 3.167 0.001
@@
Xwin-Math-7B-V1.0 0.060 0.552 0.492 3.040 0.001
Meta-Llama-3-8B 0.054 0.502 0.448 2.734 0.003
phi-1.5 0.051 0.324 0.274 2.814 0.002
Phind-CodeLlama-34B-v2 0.049 0.419 0.370 2.531 0.006
CodeLlama-34b-Instruct-hf 0.045 0.426 0.381 2.338 0.010
Phi-3-medium-128k-instruct 0.044 0.869 0.825 3.103 0.001
CodeLlama-13b-Python-hf 0.044 0.223 0.179 2.759 0.003
gemma-7b 0.043 0.519 0.476 2.198 0.014
Phi-3-mini-4k-instruct 0.040 0.788 0.748 2.385 0.009
Yi-34B-Chat 0.035 0.685 0.650 1.883 0.030
mistral-medium-latest 0.035 0.790 0.755 2.104 0.018
Mixtral-8x7B-v0.1 0.035 0.591 0.557 1.771 0.038
Xwin-Math-70B-V1.0 0.034 0.806 0.772 2.107 0.018
Mixtral-8x7B-Instruct-v0.1 0.030 0.660 0.630 1.588 0.056
Mistral-7B-v0.1 0.027 0.391 0.364 1.421 0.078
Mixtral-8x22B-Instruct-v0.1 0.026 0.872 0.846 1.913 0.028
CodeLlama-70b-Instruct-hf 0.026 0.513 0.486 1.323 0.093
Llama-2-7b-hf 0.025 0.141 0.116 1.892 0.029
Mistral-7B-Instruct-v0.1 0.025 0.353 0.329 1.309 0.095
CodeLlama-70b-hf 0.024 0.478 0.454 1.221 0.111
gemma-7b-it 0.023 0.325 0.302 1.247 0.106
mistral-small-latest 0.022 0.790 0.768 1.343 0.090
CodeLlama-13b-hf 0.021 0.236 0.215 1.247 0.106
Phi-3-medium-4k-instruct 0.020 0.874 0.854 1.519 0.064
Mixtral-8x22B-v0.1 0.020 0.767 0.748 1.138 0.127
CodeLlama-34b-hf 0.017 0.354 0.337 0.919 0.179
gemma-2b 0.015 0.185 0.170 0.966 0.167
Meta-Llama-3-70B-Instruct 0.014 0.914 0.900 1.251 0.105
CodeLlama-7b-Python-hf 0.013 0.131 0.118 1.040 0.149
dbrx-base 0.012 0.731 0.719 0.707 0.240
pythia-12b 0.011 0.036 0.025 1.701 0.044
Phi-3-mini-128k-instruct 0.011 0.757 0.746 0.645 0.260
Meta-Llama-3-70B 0.011 0.817 0.806 0.707 0.240
CodeLlama-34b-Python-hf 0.010 0.312 0.301 0.549 0.291
gpt-3.5-turbo 0.009 0.760 0.750 0.546 0.293
Mistral-7B-Instruct-v0.2 0.009 0.428 0.419 0.469 0.319
claude-3-haiku-20240307 0.009 0.785 0.776 0.532 0.298
Llama-2-70b-hf 0.008 0.552 0.544 0.445 0.328
CodeLlama-7b-Instruct-hf 0.007 0.177 0.169 0.472 0.319
gemini-1.5-pro-preview-0514 0.006 0.895 0.890 0.472 0.318
gemini-1.5-pro-preview-0409 0.005 0.897 0.892 0.403 0.343
CodeLlama-13b-Instruct-hf 0.005 0.267 0.262 0.257 0.399
dbrx-instruct 0.004 0.730 0.726 0.211 0.417
gpt-4-turbo 0.003 0.898 0.895 0.270 0.394
gpt2-xl 0.002 0.009 0.007 0.778 0.218
gpt-4o 0.002 0.931 0.929 0.219 0.413
gemini-pro -0.001 0.792 0.793 -0.081 0.532
mistral-large-latest -0.001 0.853 0.854 -0.049 0.519
gemma-2b-it -0.001 0.111 0.112 -0.106 0.542
claude-2.1 -0.004 0.887 0.891 -0.336 0.632
CodeLlama-7b-hf -0.007 0.126 0.133 -0.525 0.700
Llama-2-13b-hf -0.011 0.236 0.246 -0.629 0.735
gpt-4 -0.012 0.911 0.923 -1.161 0.877
claude-3-sonnet-20240229 -0.016 0.719 0.735 -0.894 0.814
claude-3-opus-20240229 -0.022 0.802 0.824 -1.421 0.922
deepseek-math-7b-rl -0.031 0.187 0.217 -1.963 0.975
gemini-1.5-flash-preview-0514 -0.038 0.797 0.835 -2.507 0.994
@@
math-shepherd-mistral-7b-rl 0.074 0.820 0.746 4.504 0.000
deepseek-math-7b-rl 0.064 0.760 0.696 3.672 0.000
Yi-6B-Chat 0.058 0.426 0.368 2.964 0.002
CodeLlama-34b-Python-hf 0.056 0.337 0.280 3.059 0.001
command 0.051 0.457 0.407 2.596 0.005
phi-1.5 0.050 0.321 0.271 2.786 0.003
Xwin-Math-13B-V1.0 0.042 0.662 0.620 2.212 0.013
CodeLlama-70b-Instruct-hf 0.040 0.529 0.489 2.047 0.020
CodeLlama-70b-hf 0.037 0.517 0.480 1.878 0.030
gemma-7b 0.036 0.568 0.532 1.826 0.034
gemini-1.5-pro-preview-0409 0.035 0.908 0.873 2.883 0.002
Phi-3-mini-128k-instruct 0.035 0.818 0.783 2.211 0.014
phi-2 0.034 0.552 0.518 1.744 0.041
Xwin-Math-7B-V1.0 0.033 0.530 0.497 1.680 0.046
CodeLlama-7b-hf 0.027 0.123 0.095 2.242 0.012
Mistral-7B-Instruct-v0.2 0.027 0.437 0.410 1.389 0.082
dbrx-base 0.024 0.712 0.688 1.322 0.093
gemma-7b-it 0.023 0.254 0.231 1.394 0.082
Meta-Llama-3-8B-Instruct 0.023 0.774 0.751 1.362 0.087
CodeLlama-7b-Instruct-hf 0.022 0.187 0.165 1.493 0.068
Yi-34B-Chat 0.022 0.679 0.656 1.170 0.121
mistral-small-latest 0.021 0.782 0.761 1.305 0.096
CodeLlama-13b-Python-hf 0.021 0.218 0.197 1.299 0.097
CodeLlama-34b-hf 0.020 0.330 0.310 1.097 0.136
pythia-12b 0.019 0.049 0.030 2.553 0.005
mistral-medium-latest 0.019 0.789 0.770 1.152 0.125
Phi-3-medium-4k-instruct 0.018 0.901 0.883 1.428 0.077
dbrx-instruct 0.016 0.713 0.697 0.924 0.178
Mixtral-8x7B-Instruct-v0.1 0.016 0.679 0.662 0.870 0.192
gemma-2b 0.016 0.194 0.178 1.020 0.154
Llama-2-7b-hf 0.014 0.142 0.128 1.021 0.154
Phind-CodeLlama-34B-v2 0.014 0.398 0.384 0.728 0.233
Mixtral-8x7B-v0.1 0.013 0.614 0.601 0.690 0.245
Meta-Llama-3-8B 0.013 0.547 0.534 0.660 0.255
gemini-pro 0.012 0.688 0.676 0.677 0.249
Mistral-7B-v0.1 0.011 0.431 0.420 0.583 0.280
Meta-Llama-3-70B-Instruct 0.008 0.907 0.899 0.714 0.238
Mixtral-8x22B-Instruct-v0.1 0.008 0.890 0.882 0.612 0.270
Phi-3-mini-4k-instruct 0.007 0.807 0.800 0.474 0.318
claude-3-haiku-20240307 0.006 0.792 0.785 0.416 0.339
Llama-2-13b-hf 0.005 0.281 0.276 0.298 0.383
Xwin-Math-70B-V1.0 0.005 0.808 0.803 0.319 0.375
gpt2-xl 0.004 0.006 0.002 1.422 0.078
Mixtral-8x22B-v0.1 0.002 0.808 0.807 0.115 0.454
gemini-1.5-flash-preview-0514 0.001 0.810 0.808 0.110 0.456
CodeLlama-7b-Python-hf 0.001 0.119 0.118 0.112 0.455
CodeLlama-13b-Instruct-hf -0.000 0.284 0.285 -0.028 0.511
gemma-2b-it -0.000 0.101 0.101 -0.064 0.526
CodeLlama-34b-Instruct-hf -0.002 0.403 0.404 -0.073 0.529
CodeLlama-13b-hf -0.004 0.213 0.217 -0.232 0.592
Phi-3-medium-128k-instruct -0.005 0.870 0.876 -0.368 0.644
claude-3-opus-20240229 -0.006 0.830 0.836 -0.396 0.654
claude-2.1 -0.006 0.836 0.842 -0.425 0.665
gpt-4 -0.008 0.919 0.927 -0.790 0.785
gpt-4-turbo -0.011 0.847 0.858 -0.825 0.795
Mistral-7B-Instruct-v0.1 -0.011 0.340 0.352 -0.617 0.731
gpt-4o -0.012 0.913 0.925 -1.188 0.882
Llama-2-70b-hf -0.013 0.572 0.585 -0.636 0.738
gemini-1.5-pro-preview-0514 -0.014 0.802 0.816 -0.894 0.814
mistral-large-latest -0.017 0.854 0.871 -1.228 0.890
gpt-3.5-turbo -0.017 0.742 0.759 -0.994 0.840
claude-3-sonnet-20240229 -0.024 0.713 0.737 -1.326 0.908
Meta-Llama-3-70B -0.034 0.815 0.849 -2.287 0.989
Appendix G 50 Examples from GSM1k
A previous version of this paper mistakenly included some questions from a nonfinal version of GSM1k. A corrected table is below.
@@
4 Colin is playing dice with his friend Eoin and needs some help keeping track of his score. He begins with 5 points and wins 6 points in the first round. In the second round, he won twice as many points as he won in the first round. In the third round, he had a fantastic roll and was able to triple his total point count! How many points did Colin end the game with? 69
@@
21 Frank sneaks out before his break at 3:20 pm and gets back at 4:05. If his break was only supposed to be half an hour, for how much longer did Frank sneak out? 15
22 Janet wants to listen to 20 music albums by the end of the week. If she just finished her twelfth album and today is Thursday, how many albums per day would she have to listen to by Saturday? 4
@@
43 Johnny has 8 green balls. He has five fewer than twice that in red balls. How many total balls does Johnny have? 19
@@
47 The highest temperature ever recorded on Earth was 136 degrees Fahrenheit and the coldest temperature ever measured was -126 degrees Fahrenheit. If the average temperature of Earth is 59, what would be the difference between the average temperature on Earth and the average given the two extremes? 54
48 Maria was shopping for the perfect prom dress. She found a red one that cost \$250 but was on sale for 20% off. Sales tax is 5%. Her grandmother gave her \$300 to pay for her dress and dinner. After Maria purchased the red dress, how much did she have left, in dollars, to pay for dinner? 90
49 Mrs. Watson, a high school Spanish teacher, is required to input 2 grades a week per student per her school’s grading policy. Mrs. Watson has 6 classes in total. Her 1st-period class has 32 students, 2nd-period has 28 students, 3rd-period has 41 students, 4th-period has 23 students, 5th-period has 18 students, and her 6th-period class has 33 students. How many grades does Mrs. Watson need to input each week to remain compliant with the school’s grading policy? 350
50 Mary sells 4 bags of pears where each bag contains 3 giant pears and 4 small pears. Each giant pear is sold at \$5 and each small pear is sold at \$2. Mary also sells 2 bags of cherries where each bag contains 5 pounds of cherries. The cherries are sold at \$8 per pound. How much in total does Mary earn from selling all these fruits? 172
Appendix H Bar Chart of Performance Gaps Between GSM8k and GSM1k Across All Model Accuracies